X-ray emission from PTT stars

Giulio Del Zanna

Summary

Department of Applied Mathematics and Theoretical Physics University of Cambridge • Coordinated XMM-Newton and ground-based spectroscopic +

photometric observations of two Post T Tauri stars.

• The observations were done in February → results are only preliminary !

Acknowledgements

I would like to acknowledge the participation, contribution/support from:

• Gordon Bromage and Hannah Leigh Worters, UCLAN, UK

• Helen Mason, University of Cambridge UK

• Carl Foley, Hilary McKay, MSSL, UK

• Alan Whiting group (CTIO, Chile), in particular Melissa Rice.

• Carlos Alberto, Germano Quast and Ramiro Delareza (Brasil).

• The XMM/SOC at Villafranca, Madrid (Spain), for their professional support. In particular, of Nora Loiseau and Pedro M. Rodriguez Pascual.

• Federico Gonzalez (CASLEO), Argentina

• Massimo Landini, Univ. of Florence Italy

Goals

• PPARC • Determine the physical characteristics of the X-ray emission (during

quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

• Determine the physical characteristics of the X-ray emission (during quiescent and flaring state) of PTT stars to help understand the evolution from the T-Tauri phase;

• relate them to the properties of the stellar photospheres;

• in particular in terms of chemical composition (study the FIP/anti-FIP effect).

• It is possible that the X-ray emission from CTTS is due to accretion (as proposed by Kastner et al. 2002 for TW Hya)

• the PTTS have strong X-ray emission, and could have extended solar-like coronae.

• This is perhaps the first XMM observation. Only one other PTTS (PZ Tel, Argiroffi et al. 2004).

• Multiwavelengthmultiwavelength

multiwavelength

The target

• 2RE J0241-53 (EUVE J0241-530 but we nicknamed Horace) is a binary of active X-ray emitting stars in the southern sky (Horologium), discovered by the ROSAT and EUVE all-sky surveys (see also Jeffries et al. 1996).

• classified as a member of an association of young PTT stars (∼ 30 Myr) by the brasilians (Torres et al. 2000).

• Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

• Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995). → kill two birds with a stone... and

double the chances of observingbig flares.

• dK7e and dM3e fast rotating stars (v sin i=12, 70 km/s). Southern star is Li-rich.

• 2RE J0241-53 (EUVE J0241-530 but we nicknamed Horace) is a binary of active X-ray emitting stars in the southern sky (Horologium), discovered by the ROSAT and EUVE all-sky surveys (see also Jeffries et al. 1996).

• classified as a member of an association of young PTT stars (∼ 30 Myr) by the brasilians (Torres et al. 2000).

• Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

• Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995). → kill two birds with a stone... and

double the chances of observingbig flares.

• dK7e and dM3e fast rotating stars (v sin i=12, 70 km/s). Southern star is Li-rich.

• 2RE J0241-53 (EUVE J0241-530 but we nicknamed Horace) is a binary of active X-ray emitting stars in the southern sky (Horologium), discovered by the ROSAT and EUVE all-sky surveys (see also Jeffries et al. 1996).

• classified as a member of an association of young PTT stars (∼ 30 Myr) by the brasilians (Torres et al. 2000).

• Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

• Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995). → kill two birds with a stone... and

double the chances of observingbig flares.

• dK7e and dM3e fast rotating stars (v sin i=12, 70 km/s). Southern star is Li-rich.

• 2RE J0241-53 (EUVE J0241-530 but we nicknamed Horace) is a binary of active X-ray emitting stars in the southern sky (Horologium), discovered by the ROSAT and EUVE all-sky surveys (see also Jeffries et al. 1996).

• classified as a member of an association of young PTT stars (∼ 30 Myr) by the brasilians (Torres et al. 2000).

• Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

• Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995). → kill two birds with a stone... and

double the chances of observingbig flares.

• dK7e and dM3e fast rotating stars (v sin i=12, 70 km/s). Southern star is Li-rich.

• 2RE J0241-53 (EUVE J0241-530 but we nicknamed Horace) is a binary of active X-ray emitting stars in the southern sky (Horologium), discovered by the ROSAT and EUVE all-sky surveys (see also Jeffries et al. 1996).

• classified as a member of an association of young PTT stars (∼ 30 Myr) by the brasilians (Torres et al. 2000).

• Count rates in the 100 Å bandpassess were comparable to those of the brightest flare stars.

• Both stars have strong chromospheric emission and have recurrent flares (Ball and Bromage 1995). → kill two birds with a stone... and

double the chances of observingbig flares.

• dK7e and dM3e fast rotating stars (v sin i=12, 70 km/s). Southern star is Li-rich.

Pushing to the limits

• Problems in coordinating multi-wavelength observations • XMM OM fast mode windows

• angular separation of 22” – roll angle

• multi-continent, multi-cultural experience

• we need an observatory on Easter Island!

Ground-based

Two-weeks observing in South Africa. Spectroscopy

• SAAO 1.9m grating spectrograph with 1 Å resolution and ≃90 s exposures.

• Wavelength range: 3700 - 4360 Å

• 6 consecutive nights Photometry

• UBV aperture photometry with the SAAO 0.5-m telescope

• 3 nights

On the XMM night (19-20 Feb 2004)

• Argentina – CASLEO: cloudy !

• Brasil: B-band monitoring 23:00 – 00:30 UT then cloudy.

• Chile:

• CTIO low-resolution (17.2 Å FWHM) spectra at the 1.5m • CTIO CCD B and U band photometry with the 0.9m 01:17

04:40 UT

• South Africa:

• SAAO 1.9m spectra 18:32 UT – 00:04 UT

• SAAO U-band monitoring

→ ground-based only covered part of the XMM observation

SAAO spectra - northern star

Two spectra on the 7th and 9th Feb 2004:

Filters - CTIO

COSPAR 2004: X-ray emission from PTT stars – p.10/29

Filters - OM

COSPAR 2004: X-ray emission from PTT stars – p.11/29

Chromospheric variability

Consecutive 90s exposures from XMM night (19:12,19:14 UT). Notice the variation in the Ca II (H and K) and H δ.

EW for the Northern star

COSPAR 2004: X-ray emission from PTT stars – p.13/29

XMM OM

COSPAR 2004: X-ray emission from PTT stars – p.14/29

OM light curves

COSPAR 2004: X-ray emission from PTT stars – p.15/29

EPIC MOS 1 image

COSPAR 2004: X-ray emission from PTT stars – p.16/29

EPIC PN light curves

COSPAR 2004: X-ray emission from PTT stars – p.17/29

EPIC PN light curves

COSPAR 2004: X-ray emission from PTT stars – p.18/29

Flares on Northern star

COSPAR 2004: X-ray emission from PTT stars – p.19/29

Flares on Northern star

COSPAR 2004: X-ray emission from PTT stars – p.20/29

Flares on the Southern star.

COSPAR 2004: X-ray emission from PTT stars – p.21/29

Flares on the Southern star.

COSPAR 2004: X-ray emission from PTT stars – p.22/29

EPIC MOS 1 spectrum of the Southern star

COSPAR 2004: X-ray emission from PTT stars – p.23/29

RGS - here is why we need spectroscopy

Where is all the Iron gone ?

Gallery - (from Huenemoerder, 2002)

COSPAR 2004: X-ray emission from PTT stars – p.25/29

Horace vs. PZ Tel

COSPAR 2004: X-ray emission from PTT stars – p.26/29

Densities - Ne IX

COSPAR 2004: X-ray emission from PTT stars – p.27/29

Densities - O VII

COSPAR 2004: X-ray emission from PTT stars – p.28/29

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).

• This is one of the few (two ?) X-ray observations of PTT stars.

• The XMM/OM fast-mode windows worked very well but data are missing.

• Both stars confirmed to be active, with variability and flares.

• The coronae are mostly abundant in Ne,O.

• Compared to other young stars and PZ Tel, these stars have no ‘hot’ emission (more like CTTS). We need more observations.

• They were in a remarkably quiescent state (stellar cycles?).

• Ground-based SAAO and CTIO observations are nicely complementary but missed all the action !

• We confirm Linsky’s law (the most interesting events occurr at the end of the observation).